# Quantitative estimates for singularity for conjugate equations driven by linear fractional transformations

### v3

@@ 3.1. Proofs
$\liminf_{n\to\infty}\frac{|\{1\leq j\leq n:\xi_{j}(z)\in I\}|}{n}\geq(1-\eta)c_{I},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
The assertion follows if $\eta\to+0$. ∎
@@
(ii) If $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$, then, $\alpha=-1,\beta=+\infty$.
(iii) If $a_{0}=1$ and $b_{N-1}+c_{N-1}>0$, then, $\alpha>-1,\beta=+\infty$.
(iv) If $a_{0}<1$ and $b_{N-1}+c_{N-1}=0$, then, $\alpha=-1,\beta<+\infty$.
(v) If $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$, then, $\alpha>-1,\beta<+\infty$.
(vi) If $a_{0}=1$, then, $\textup{Fix}(H_{0})=\{+\infty\}$ and $\mathfrak{p}_{0}=(1,0,\cdots,0)\in\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$.
(vii) If $b_{N-1}+c_{N-1}=0$, then, $\textup{Fix}(H_{N-1})=\{-1\}$ and $\mathfrak{p}_{N-1}=(0,\cdots,0,1)\in\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$.
@@
Case 1. We first deal with the case that $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$.
In this case, we do not need to assume ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
@@
$10<\inf_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)\leq\sup_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)<$ 1 .
Fix ${\bf q}\in\overline{\mathcal{P}_{N}}$. Let $\widetilde{c}:=\inf_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)$. Then, $c_{\{\mathscr{A}({\bf q})\}}\geq\widetilde{c}>0$ and by Lemma 3.13 for $I=\{\mathscr{A}({\bf q})\}$,
$\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}{\bf 1}_{\{\xi_{k}=\mathscr{A}({\bf q})\}}(z)\geq\widetilde{c},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
The number $i_{1}\in\mathcal{I}_{N}$ in Lemma 3.11 will be chosen according to the position of ${\bf q}$. By applying Lemma 3.11 to $i_{1}=\mathscr{A}({\bf q})$ and $\epsilon_{1}=\widetilde{\epsilon_{1}}$ where $\widetilde{\epsilon_{1}}$ is the constant in Lemma 3.12, we see that if $X_{i+1}(z)=\mathscr{A}({\bf q})$, then,
$s_{N}\left(p_{i}(0;z)|{\bf q}\right)+s_{N}\left(p_{i+1}(0;z)|{\bf q}\right)\leq V(\widetilde{\epsilon_{1}})<0.$
Now the assertion follows if we let $\epsilon_{2}:=\dfrac{\widetilde{c}\ V(\widetilde{\epsilon_{1}})}{2}$. ∎
Case 2. We now deal with the case that $a_{0}=1$ or $b_{N-1}+c_{N-1}=0$.
If $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$, then, $c_{\{0\}}=c_{\{N-1\}}=0$, so we consider $c_{\{0,N-1\}}$. In the following, we do not use Lemma 3.12 or the assumption that |$|\{\mathfrak{p}_{i}\}_{i}|\geq 2$. On the other hand, we assume ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
@@
Case 2.1. We first consider the case that $a_{0}=1$ and $b_{N-1}+c_{N-1}>0$.
By Lemma 3.14, $\alpha>-1$ and $\beta=+\infty$, $H_{0}(y)=y+c_{0}$, and $\textup{Fix}(H_{0})=\{+\infty\}$. $\mathfrak{p}_{0}=(1,0,\cdots,0)\notin\mathcal{P}_{N}$. On the other hand, $\mathfrak{p}_{i}\in\mathcal{P}_{N}$ for $i\geq 1$.
Since $G_{0}(+\infty)=G_{0}(H_{0}(+\infty))=1$ and $q_{0}<1-\epsilon/2$,
$\inf_{y\in Y}W_{0}({\bf q},y)>c_{0}(\epsilon/2)>0,$
where the lower bound $c_{0}(\epsilon/2)$ is uniform with respect to ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
Since $\alpha>-1$, $\inf_{y\in Y}G_{0}(y)\geq c>0$. By using Lemma 3.11 and Lemma 3.13 for $I=\{0\}$, we have the assertion.
Case 2.2. We secondly consider the case that $a_{0}<1$ and $b_{N-1}+c_{N-1}=0$.
We recall that $\alpha=-1$ and $\beta<+\infty$, By Lemma 3.14, $H_{N-1}(y)=\dfrac{(1-2b_{N-1})y+b_{N-1}}{1-b_{N-1}}$, and $\textup{Fix}(H_{N-1})=\{-1\}$. $\mathfrak{p}_{N-1}=(0,\cdots,0,1)\notin\mathcal{P}_{N}$. On the other hand, $\mathfrak{p}_{i}\in\mathcal{P}_{N}$ for $i\leq N-2$.
@@
$\inf_{y\in Y}W_{N-1}({\bf q},y)>c_{N-1}(\epsilon/2)>0,$
where the lower bound $c_{N-1}(\epsilon)$ is uniform with respect to ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$. Since $\beta<+\infty$, $\inf_{y\in Y}G_{N-1}(y)\geq c>0$. The rest of the argument is the same as in the above case.
Case 2.3. We finally consider the case that $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$.
Since $\textup{Fix}(H_{0})=\{+\infty\}$, $\textup{Fix}(H_{N-1})=\{-1\}$, $q_{0}<1-\epsilon/2$ and $q_{N-1}<1-\epsilon/2$, it holds that
$\inf_{y\in Y}W_{0}({\bf q},y)>c_{0}(\epsilon/2)>0$
@@
Now the proof of Theorem 3.2 completes.
@@
(ii) Assume that $\{g_{i}\}_{i}$ are affine functions such that $g_{i}(x)=b_{i}x+(b_{0}+\cdots+b_{i-1})$. Since $Y=\{0\}$, there is no need to introduce $H_{i}$, and $W_{i}({\bf q},y)=\max_{k}|G_{k}(y)-q_{k}|=\max_{k}|b_{k}-q_{k}|$. However, $\mathscr{p}_{0}=\cdots=\mathscr{p}_{N-1}=(b_{i})_{i}$, and hence this is not in the scope of Theorem 3.2. This case is considered in Theorem 3.3.
@@
By (3.6), if we take $\delta>0$ such that
$\frac{\log N}{\log(N/(1+\delta))}\leq 1+\frac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)},$ (3.9)
then, $s\leq 1+\dfrac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)}$.
Assume that $s\leq 1+\dfrac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)}$. Since $\min_{i}r_{i}\geq\epsilon$, $\sum_{i}|r_{i}-r_{i}^{s}|\leq\epsilon^{1-s}-1\leq\epsilon/2$. Hence, $\left\|{\bf\widetilde{q}}-\mathfrak{p}_{0}\right\|\geq\epsilon/2$, where we let $\widetilde{q}_{i}:=r_{i}^{s}$ for each $i\in\mathcal{I}_{N}$ and ${\bf\widetilde{q}}:=\left(\widetilde{q}_{i}\right)_{i\in\mathcal{I}_{N}}$. Since $\epsilon\leq 1/2$, $\widetilde{q}_{i}\geq\epsilon^{s}\geq\epsilon/2$ for each $i\in\mathcal{I}_{N}$. Hence, ${\bf\widetilde{q}}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
@@
$\epsilon_{3}:=\inf\left\{W_{i}({\bf q},y)\middle|i\in\mathcal{I}_{N},y\in Y,{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)},\left\|{\bf q}-\mathfrak{p}_{0}\right\|\geq\epsilon/2\right\}>0.$
By Lemma 3.11, we see that for every $j\geq 0$, every $z\in(\mathcal{I}_{N})^{\mathbb{N}}$, and every ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$ with $\left\|{\bf q}-\mathfrak{p}_{0}\right\|\geq\epsilon/2$,
@@
By Lemma 3.9, we can let
$V_{\epsilon}(\epsilon_{3}):=\max\left\{V\left({\bf q};\epsilon_{3}\right)\middle|{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)},\left\|{\bf q}-\mathfrak{p}_{0}\right\|\geq\epsilon/2\right\}<0.$
@@
$\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq-\frac{V_{\epsilon}(\epsilon_{3})}{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
Assume that $\delta>0$ satisfies not only (3.9) but also (3.5) for $\epsilon_{2}:=V_{\epsilon}(\epsilon_{3})/2$. Then, by Propositions 3.6 and 3.7, $\dim_{H}\mu_{\varphi}<1$.
@@
(i) We do not need to use Lemma 3.13. For every small $\epsilon>0$ and every $i\in\mathcal{I}_{N}$,
@@
by Lemma 3.10 (ii).
(ii) The case that both of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are affine is easily handled by the dimension formula. However, if some of the functions $\{f_{i},g_{i}\}_{i}$ are non-affine, then, it is difficult to judge the singularity by a generalized dimension formula in [9].
4. Example
We give three examples that $\{g_{i}\}_{i}$ is an LF system and $\{f_{i}\}_{i}$ is non-affine and the solution $\varphi$ is singular. The result depends on the quantitative estimates in Section 3.
We first deal with the case that $N=2$. Then,

### v4

@@ 3.1. Proofs
$\liminf_{n\to\infty}\frac{|\{1\leq j\leq n\colon\xi_{j}(z)\in I\}|}{n}\geq(1-\eta)c_{I},\ \textup{ $\displaystyle\nu_{0}$-a.s.$\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
The assertion follows for $\eta\to+0$. ∎
@@
(ii) If $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$, then $\alpha=-1,\beta=+\infty$.
(iii) If $a_{0}=1$ and $b_{N-1}+c_{N-1}>0$, then $\alpha>-1,\beta=+\infty$.
(iv) If $a_{0}<1$ and $b_{N-1}+c_{N-1}=0$, then $\alpha=-1,\beta<+\infty$.
(v) If $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$, then $\alpha>-1,\beta<+\infty$.
(vi) If $a_{0}=1$, then $\textup{Fix}(H_{0})=\{+\infty\}$ and $\mathfrak{p}_{0}=(1,0,\ldots,0)\in\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$.
(vii) If $b_{N-1}+c_{N-1}=0$, then $\textup{Fix}(H_{N-1})=\{-1\}$ and $\mathfrak{p}_{N-1}=(0,\ldots,0,1)\in\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$.
@@
Case 1. We first deal with the case where $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$.
@@
$0<\inf_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)\leq\sup_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)<1.$
Fix ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$. Let $\displaystyle\widetilde{c}\coloneqq\inf_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)$. Then,
$c_{\{\mathscr{A}({\bf q})\}}=\inf_{y\in Y}G_{\mathscr{A}({\bf q})}(y)\geq\widetilde{c}>0,$
and by Lemma 3.13 for $I=\{\mathscr{A}({\bf q})\}$,
$\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}{\bf 1}_{\{\xi_{k}=\mathscr{A}({\bf q})\}}(z)\geq\widetilde{c},\ \textup{ $\displaystyle\nu_{0}$-a.s. $\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
The number $i_{1}\in\mathcal{I}_{N}$ in Lemma 3.11 will be chosen according to the position of ${\bf q}$. By applying Lemma 3.11 to $i_{1}=\mathscr{A}({\bf q})$ and $\epsilon_{1}=\widetilde{\epsilon_{1}}$ where $\widetilde{\epsilon_{1}}$ is the constant in Lemma 3.12, we see that if $\xi_{i+1}(z)=\mathscr{A}({\bf q})$, then
$s_{N}\left(p_{i}(0;z)|{\bf q}\right)+s_{N}\left(p_{i+1}(0;z)|{\bf q}\right)\leq V_{\epsilon,1}(\widetilde{\epsilon_{1}})<0.$
Now the assertion follows if we let $\epsilon_{2}\coloneqq-\dfrac{\widetilde{c}\ V_{\epsilon,1}(\widetilde{\epsilon_{1}})}{2}>0$. ∎
Case 2. We now deal with the case where $a_{0}=1$ or $b_{N-1}+c_{N-1}=0$.
If $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$, then $c_{\{0\}}=c_{\{N-1\}}=0$, so we consider $c_{\{0,N-1\}}$. In the following, we use neither Lemma 3.12 nor the assumption that $\left|\left\{\mathfrak{p}_{i}\right\}_{i}\right|\geq 2$.
@@
We consider the following three cases.
Case 2.1. We consider the case where $a_{0}=1$ and $b_{N-1}+c_{N-1}>0$.
By Lemma 3.14, we see that $\alpha>-1$ and $\beta=+\infty$. It holds that $H_{0}(y)=y+c_{0}$, and hence $\textup{Fix}(H_{0})=\{+\infty\}$. Hence, $\mathfrak{p}_{0}={\bf G}(\textup{Fix}(H_{0}))=(1,0,\ldots,0)\notin\mathcal{P}_{N}$. On the other hand, $\mathfrak{p}_{i}\in\mathcal{P}_{N}$ for $i\geq 1$.
Since $G_{0}(+\infty)=G_{0}(H_{0}(+\infty))=1$ and $q_{0}\leq 1-\epsilon/2$,
$\inf_{y\in Y}W_{0}({\bf q},y)\geq\widetilde{c_{0}}(\epsilon/2),$
where the positive lower bound $\widetilde{c_{0}}(\epsilon/2)$ is uniform with respect to ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
Since $\alpha>-1$, it holds that $\displaystyle\inf_{y\in Y}G_{0}(y)\geq c$ for some $c>0$. By using Lemma 3.11 and Lemma 3.13 for $I=\{0\}$, we have the assertion.
Case 2.2. We consider the case where $a_{0}<1$ and $b_{N-1}+c_{N-1}=0$.
By Lemma 3.14, we see that $\alpha=-1$ and $\beta<+\infty$. It holds that $H_{N-1}(y)=\dfrac{(1-2b_{N-1})y+b_{N-1}}{1-b_{N-1}}$, and hence $\textup{Fix}(H_{N-1})=\{-1\}$. Hence, $\mathfrak{p}_{N-1}={\bf G}(\textup{Fix}(H_{N-1}))=(0,\ldots,0,1)\notin\mathcal{P}_{N}$. On the other hand, $\mathfrak{p}_{i}\in\mathcal{P}_{N}$ for $i\leq N-2$.
@@
$\inf_{y\in Y}W_{N-1}({\bf q},y)\geq\widetilde{c_{N-1}}(\epsilon/2),$
where the positive lower bound $\widetilde{c_{N-1}}(\epsilon)$ is uniform with respect to ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$. Since $\beta<+\infty$, it holds that $\displaystyle\inf_{y\in Y}G_{N-1}(y)\geq c$ for some $c>0$. The rest of the argument is the same as in the above case.
Case 2.3. We consider the case where $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$.
Since $\textup{Fix}(H_{0})=\{+\infty\}$, $\textup{Fix}(H_{N-1})=\{-1\}$, $q_{0}\leq 1-\epsilon/2$ and $q_{N-1}\leq 1-\epsilon/2$, it holds that
$\inf_{y\in Y}W_{0}({\bf q},y)>\widetilde{c_{0}}(\epsilon/2)>0$
@@
Now the proof of Theorem 3.2 is completed.
@@
(ii) Assume that $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are affine maps such that $g_{i}(x)=b_{i}x+(b_{0}+\cdots+b_{i-1})$. Since $Y=\{0\}$, there is no need to introduce $H_{i}$, and $W_{i}({\bf q},y)=\max_{k}|G_{k}(y)-q_{k}|=\max_{k}|b_{k}-q_{k}|$. However, it holds that $\mathfrak{p}_{0}=\cdots=\mathfrak{p}_{N-1}=(b_{i})_{i}$, and hence this is not in the scope of Theorem 3.2. This case is considered in Theorem 3.3.
@@
By Lemma 3.4, if we take $\delta>0$ such that
(3.10) $\frac{\log N}{\log(N/(1+\delta))}\leq 1+\frac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)},$
then $\displaystyle s\leq 1+\dfrac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)}$.
Assume that $\displaystyle s\leq 1+\dfrac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)}$. Since $\displaystyle\min_{i\in\mathcal{I}_{N}}r_{i}\geq\epsilon$, $\displaystyle\sum_{i\in\mathcal{I}_{N}}|r_{i}-r_{i}^{s}|\leq\epsilon^{1-s}-1\leq\epsilon/2$. Hence, $\left\|{\bf\widetilde{q}}-\mathfrak{p}_{0}\right\|_{1}\geq\epsilon/2$, where we let $\widetilde{q}_{i}\coloneqq r_{i}^{s}$ for each $i\in\mathcal{I}_{N}$ and ${\bf\widetilde{q}}\coloneqq\left(\widetilde{q}_{i}\right)_{i\in\mathcal{I}_{N}}$. Since $\epsilon\in(0,1)$, $\widetilde{q}_{i}\geq\epsilon^{s}\geq\dfrac{\epsilon}{1+\epsilon/2}\geq\epsilon/2$ for each $i\in\mathcal{I}_{N}$. Hence, ${\bf\widetilde{q}}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
@@
$\epsilon_{3}\coloneqq\inf\left\{W_{i}({\bf q},y)\middle|i\in\mathcal{I}_{N},y\in Y,{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)},\left\|{\bf q}-\mathfrak{p}_{0}\right\|_{1}\geq\epsilon/2\right\}>0.$
By Lemma 3.11, we see that for every $j\geq 0$, every $z\in(\mathcal{I}_{N})^{\mathbb{N}}$, and every ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$ with $\left\|{\bf q}-\mathfrak{p}_{0}\right\|_{1}\geq\epsilon/2$,
@@
By Lemma 3.9, we see that
$V_{\epsilon,2}(\epsilon_{3})\coloneqq\max\left\{V\left({\bf q};\epsilon_{3}\right)\middle|{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)},\left\|{\bf q}-\mathfrak{p}_{0}\right\|_{1}\geq\epsilon/2\right\}<0.$
@@
$\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq\frac{V_{\epsilon,2}(\epsilon_{3})}{2},\ \textup{ $\displaystyle\nu_{0}$-a.s. $\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
Assume that $\delta>0$ satisfies not only (3.10) but also (3.6) for $\epsilon_{2}\coloneqq V_{\epsilon}(\epsilon_{3})/2$. Then, by Propositions 3.6 and 3.7, $\dim_{H}\mu_{\varphi}<1$.
@@
(i) If $\mathfrak{p}_{0}=\cdots=\mathfrak{p}_{N-1}$, then $\mathfrak{p}_{0}\in\mathcal{P}_{N}$. We show this. Since $G_{0}(y)=\dfrac{b_{1}(y+1)}{b_{1}y+1}$ and $0<b_{1}<1$, it is injective on $Y$. Hence, $\textup{Fix}(H_{0})=\cdots=\textup{Fix}(H_{N-1})$. Hence, $\displaystyle\frac{c_{0}}{1-a_{0}}=\frac{a_{1}-1+\sqrt{(1-a_{1})^{2}+4b_{1}c_{1}}}{2b_{1}}<\infty$. By Definition 2.8 and the assumption that $d_{0}=1$, we see that $a_{0}\in(0,1)$ and $\dfrac{a_{0}}{c_{0}+1}=b_{1}\in(0,1)$. Hence, $\displaystyle\frac{c_{0}}{1-a_{0}}>-1$. Recall that $\displaystyle G_{i}(y)=\frac{(b_{i+1}-b_{i})(y+1)}{(b_{i}y+1)(b_{i+1}y+1)},i\in\mathcal{I}_{N}$. Since $0=b_{0}<b_{1}<\cdots<b_{N-1}<b_{N}=1$, $G_{i}(y)>0$ for every $y\in(-1,+\infty)$ and every $i\in\mathcal{I}_{N}$. Hence, $\displaystyle\mathfrak{p}_{0}=\left(G_{i}\left(\frac{c_{0}}{1-a_{0}}\right)\right)_{i\in\mathcal{I}_{N}}\in\mathcal{P}_{N}$.
(ii) In the proof of Theorem 3.3, we do not need to use Lemma 3.13, since we have the estimate that $\epsilon_{3}>0$. We remark that for every small $\epsilon>0$ and every $i\in\mathcal{I}_{N}$,
@@
because, by (i) above, we can let ${\bf q}={\bf G}\left(\textup{Fix}(H_{i})\right)=\mathfrak{p}_{i}=\mathfrak{p}_{j},j\in\mathcal{I}_{N}$ and $y=\textup{Fix}(H_{i})$.
(iii) The case where both $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are affine maps is easily handled by the dimension formula. However, if some of the functions $\{f_{i},g_{i}\}_{i}$ are non-affine maps, then it is difficult to judge the singularity by the generalized dimension formula in [11, Corollary 3.5].
4. Examples
We give three examples assuming that $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ is an LF system and $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ are non-affine maps and the solution $\varphi$ is singular. The result depends on the quantitative estimates in Section 3. We assume that $d_{i}=1$ for every $i\in\mathcal{I}_{N}$.
We first deal with the case where $N=2$. Then,